# Enforcing Paraphrase Generation via Controllable Latent Diffusion

### v1

@@ 4.1 Setups
LDP is implemented by N=12 layers of Transformer blocks with 12 attention heads, where we adopt time step embedding in the same way as the position embedding. The maximum sequence length is $96$. We adopt BART-base (139M) for encoder and decoder, and the diffusion dimension is $768$, the same as BART embedding dimension.
The diffusion process is defined by $T=1000$, with cosine schedule Nichol and Dhariwal (2021) for $\beta_{t}$. We train the model for up to $250$k steps with batch size $192$. The learning rate for the DPM training is $10^{-4}$, while for the controller fine-tuning is $10^{-5}$. We apply DPM-Solver++Lu et al. (2022b) with 25 steps for diffusion sampling.
We trained our model on 4$\times$ v100 and 4$\times$ Nvidia 3090 GPUs for about 100 GPU-hours. LDP has approximately 200M trainable parameters.
@@
We adopt Quora Question Pairs (QQP) [^1] https://www.kaggle.com/c/ quora-question-pairs, and Twitter-URL (Twitter) Lan et al. (2017), which is popular amongst mainstream paraphrase generations. The QQP data is mass-extracted from Quora regarding the shared utterance, while the Twitter data is annotated from social media by semantic similarity. We randomly divide both datasets into three parts: 10k test sentences, 10k validation sentences, and the remaining sentences were assigned to the training set.
@@
We first choose several mainstream paraphrase generations as baselines, including the deterministic paradigm by generic Transformer Vaswani et al. (2017), fine-tuned BART-base (BART-FT) Lewis et al. (2019), and variational paradigm by T5-GPVAE Du et al. (2022) and BART-CVAE Wang and Wan (2019). We also include state-of-the-art text generations via the diffusion model, such as DiffuSeq Gong et al. (2022) and SeqDiffuSeq Yuan et al. (2022), which are versatile end-to-end diffusion modeling. DiffuSeq applies minimum Bayes risk decoding (Kumar and Byrne, 2004, MBR) with 10 candidates, while SeqDiffuSeq implements beam search for decoding given an on-the-fly encoder-decoder framework during rounding. For all beam search in the experiments, we apply beam size as $4$.
@@
src how is black money gon na go off with no longer the use of same 500 and 1000 notes ?
origin paraphrase how does black money brought out to black money market or corruption?
@@
enforced parphrase how does banning 500 and 1000 rupee notes solve black money problem?
Table 2: Enforce paraphrase semantics by input segments. We inject input segments masked by placeholder <M> of ‘black money’, ‘no longer’, and ‘1000’ via controller, which improves the paraphrase semantic.
An ideal paraphrase must achieve both generation quality and diversity. We validate text quality by reference-oriented BLEU Papineni et al. (2002) and perplexity (PPL) based on GPT2 Radford et al. (2019); The diversity is measured by intra-diversity within each generated sentence, where we adopt distinct unigram div-4 Deshpande et al. (2019). Note that reference-oriented metrics such as BLEU contradict the intuition of generation diversity, we finally adopt iBScore Dou et al. (2022) for overall metric, which measures the semantics by cosine similarity of BERT sentence embeddings, namely, BERTScore Zhang* et al. (2020), and punishes source duplication by source BLEU:
@@
where we calculate BERTScore via RoBERTa-large Liu et al. (2019).
@@ 4.2 Main Results
Table 1 presents the main results of our experiments. LDP achieves comparable or even superior performance compared to mainstream baselines. Moreover, it outperforms other diffusion counterparts on QQP and Twitter test sets. LDP achieves the best performance amongst all baselines for the QQP test set.
Specifically, LDP improves overall results (iBScore) especially diversity (div-4) compared to traditional end-to-end paradigms such as fine-tuned BART, where we consider the diffusion module significantly improves the generation diversity. On the other hand, LDP also outperforms SOTA diffusion baselines such as SeqDiffuSeq in terms of BLEU and perplexity, where we consider the LDP to implement a better bridge between the diffusion process with the discrete texts than rounding. Additionally, unlike the intuition to introduce external features, the original case output improves greatly when we enforce the generation semantics with mere a segment of source inputs as shown in Table 2.
Overall, LDP achieves the best iBScore, indicating a high-quality and diverse paraphrase generation. As shown in Table 3, we generate several paraphrases by different latent sampling for generations. Though BART-CVAE and DiffuSeq are SOTA generators for diversity, they still yield relatively resemble paraphrases. LDP, on the other hand, yields better and more diverse paraphrases.
@@
DinoiserYe et al. (2023) 26.07 -
Diff-GlatQian et al. (2022) 29.86 -
SeqDiffuSeqYuan et al. (2022) 24.32 -
DiffuSeqGong et al. (2022) 24.13 83.65
@@
We additionally include more recent text diffusion generators for comparison for QQP tests. Table 4 shows that LDP achieves state-of-the-art BLEU and BERTScore amongst the diffusion baselines.
@@ 5.1 Inference Efficiency
Timelapse (s) Acceleration
Transformer 235 206.3$\times$
BART-FT 238 203.7$\times$
T5-GPVAE 3909 12.4$\times$
BART-CVAE 252 192.4$\times$
DiffuSeq(DDIM 2000) 48480 1$\times$
DiffuSeq(DDIM 500) 11530 4.2$\times$
SeqDiffuSeq(DDIM 2000) 13851 3.5$\times$
LDP(ours) 290 167.2$\times$
Table 5: Inference overhead on QQP validation set (seconds), where we set the DiffuSeq as the efficiency baseline.
By eliminating the rounding process in text diffusion models, LDP achieves improved efficiency. We compare several baseline efficiencies under their best-generation performance. As shown in Table 5, LDP is $167.2$ times faster than DiffuSeq and $50$ times faster than SeqDiffuSeq, where the diffusion overhead takes up only $18\%$ of the total timelapse. Consequently, our approach achieves a similar efficiency compared to that of the autoregressive baselines with only encoder-decoder overheads.
@@ 5.2 Domain Adaptation by Controller
+ES 36.46 14.65 (+4.49)
Table 6: Domain adaptation by controller
The semantic controller is also apt for domain adaptation. We additionally adopt ChatGPT-augmented paraphrase dataset Vladimir Vorobev (2023)(ChatGPT-Aug) as the novel data domain, then sample $10$k sentences as the test set. ChatGPT-Aug dataset consists of paraphrases generated by ChatGPT OpenAI (2022) from an ensembled dataset including QQP, SQuAD 2.0 Rajpurkar et al. (2016) and CNN news See et al. (2017).
We first train the LDP on QQP data for $200$k steps, then fine-tune it with a controller for ChatGPT-Aug data for $40$k steps. The fine-tuning follows the same routine for controller training, where the ChatGPT-Aug data pairs are training sources and references, with additional keywords from its reference as controller inputs. Table 6 shows that adaptation by controller is viable, where the fine-tuned controller substantially improves ChatGPT-Aug test BLEU, with only minor loss for the original test domain. Note that, the performance of the original data domain is retained by inference without controller inputs, which outstands from traditional data adaptation by fine-tuning.
@@ 5.3 Ablation Study for Samplers
The diffusion sampler plays an important role during inference, thus we conduct an ablation study on the diffusion sampler for comparison. Intuitively, the diffusion generation improves by more steps given the same sampler. Due to the truncation errors introduced by rounding, ODE-based samplers are not apt for text diffusion with rounding, such as DiffuSeq and SeqDiffuSeq. Thus, we additionally adopt DDIM sampler Nichol and Dhariwal (2021) with 500 and 200 sampling steps, which is also adopted for diffusion baselines. Table 7 shows that LDP still outperforms the DiffuSeq and SeqDiffuSeq given the same sampler setting. Notably, LDP still outperforms the baselines by fewer sampling steps, which were expected to be inferior.
@@ 5.4 Question Generation Ability
Considering the compatibility of LDP besides paraphrasing, we additionally validate question generation, which aims to generate the exact question by given issue descriptions. Question generation focuses on semantics upon relatively flexible utterance, thus we evaluate the BLEU and BERTScore on Quasar-T tests Dhingra et al. (2017).
As shown in Table 8, LDP with enforced semantics also achieves a comparable BLEU score with SOTA, TESS Mahabadi et al. (2023), with much improved BERTScore.
@@ 6 Related Work
Paraphrase generation is commonly modeled by end-to-end sequence generation, such as fine-tuned Transformers Vaswani et al. (2017); Lewis et al. (2019); Yang et al. (2019). However, such deterministic generation fails to ensure diversity, thus some researchers Xie et al. (2022); Sancheti et al. (2022); Vijayakumar et al. (2016); Fan et al. (2018) introduce external features to amend. Others turn to the variational generation for diversity Bowman et al. (2015). To amend the inferior generation by latent representation, some researchers further introduce pre-trained text encoder and decoder for generation Du et al. (2022); Wang and Wan (2019).
On the other hand, the diffusion model has been a popular Markov variational generation in recent years. Existing text diffusions cater to the discrete nature of language for versatile generations. DiffuSeq Gong et al. (2022) employs a single Transformer encoder and partial noising process to extend Diffusion-LM Li et al. (2022) for end-to-end text generations. They introduce discrete sampling for diffusion intervals called ‘rounding’, to ensure generation quality. However, rounding introduces truncation errors that hinder efficient skip-step diffusion sampler. BG-Diffuseq Tang et al. (2023) aims to narrow the gap between training and sampling for text diffusion via incorporating distance penalty and adaptive decay sampling. Dinoiser Ye et al. (2023), turns to mitigate truncation errors by manipulating the noise scale scheduled for text diffusion training and inference to improve its efficacy. On the other hand, SeqDiffuSeq Yuan et al. (2022) turns to a continuous diffusion modeling within an on-the-fly encoder-decoder framework. Similarly, DiffuSIA Tan et al. (2023) introduces an encoder-decoder diffusion via spiral interaction. Diff-GLAT Qian et al. (2022) incorporates residual glancing sampling, a text reconstruction via encoder-decoder, with its dropout rate as the noise scale. TESS Mahabadi et al. (2023) proposes a logit simplex space rather than embedding space for text diffusion.
@@ 7 Conclusion
We propose a novel paraphrase generation by the controllable latent diffusion model, LDP, which can further enforce semantics for paraphrase generation by harnessing only input segments instead of external features. Experiments show that LDP generates better paraphrase with superior efficiency compared to its diffusion counterparts. Further analysis shows that LDP is also applicable to other similar text generations such as question generation. Its controller is also helpful for domain adaptation.
Overall, LDP strikes a better balance between generation quality and diversity compared to mainstream baselines.
Limitations
In this work, we opt not to implement a larger Pre-trained model than the BART-base encoder and decoder, which will require more GPU memory during training. The latent diffusion is viable for the diffusion process by lower dimensions to cut down training expenses. However, the diffusion process needs to cater to the dimension of the Pre-trained latent space. Thus, this work does not explore the impact of model scaling.
We analyse our method mainly on QQP due to the restriction of the open-sourced baselines. Additionally, we are unable to explore more controller implementations other than domain adaptation due to time constraints.
Ethics Statement

### v2

@@ 4.1 Setups
We implemented LDP by N=12 layers of Transformer blocks with 12 attention heads, where we adopt time step embedding in the same way as the position embedding. The maximum sequence length is $96$. We adopt BART-base (139M) for encoder and decoder, and the diffusion dimension is $768$, the same as BART embedding dimension.
The diffusion process is defined by $T=1000$, with cosine schedule [12] for $\beta_{t}$. We train the model for up to $250$k steps with batch size $192$. The learning rate for the DPM training is $10^{-4}$, while for the controller fine-tuning is $10^{-5}$. We apply DPM-Solver++[29] with 25 steps for diffusion sampling.
We trained our model on 4$\times$ v100 and 4$\times$ Nvidia 3090 GPUs for about 100 GPU-hours. LDP has approximately 200M total parameters.
@@
We adopt Quora Question Pairs (QQP), Twitter-URL (Twitter) [37] and positive pairs from PAWS-wiki [38], which is popular amongst mainstream paraphrase generations. For QQP and Twitter data, we randomly divided both datasets into three parts: 10k test with 10k validation sentences, and the remaining were assigned to the training set.
@@
We first adopt several mainstream paraphrase generations as baselines, including the deterministic paradigm by generic Transformer [5], fine-tuned BART-base (BART-FT) [31], and variational paradigm equipped with pretrained encoder and decoder by T5-GPVAE [9] and BART-CVAE [39]. We also include SOTA text generations via the diffusion model, such as DiffuSeq [40] and SeqDiffuSeq [41], which are versatile end-to-end diffusion modeling. DiffuSeq applies minimum Bayes risk decoding [42, MBR] with 10 candidates, while SeqDiffuSeq implements beam search for decoding given an on-the-fly enc-dec framework during rounding. For all beam search in the experiments, we apply beam size as $4$. Last but not least, we adopt the latest open-source LLMs for comparison, Llama2-7b [43] and Llama3.1-8b [44] [^1] Following reviewers’ recommendations, we incorporate LLMs into our comparison, which were introduced a year after our initial draft., which are supervised fine-tuned via LoRA [45] given the same training data on the same device. We adopt $rank$ and $lora\_\alpha$ by $(64,128)$ for one-epoch LoRA tuning.
@@
src how is black money gon na go off with no longer the use of the same 500 and 1000 notes ?
original paraphrase how does black money brought out to black money market or corruption?
@@
Guided paraphrase how does banning 500 and 1000 rupee notes solve the black money problem?
Table 2: Guide paraphrase semantics by input segments. One can highlight the input segments to guide the paraphrase. We inject input segments masked by placeholder <M> of ‘black money’, ‘no longer’, and ‘1000’ via controller, which improves the paraphrase.
We first cater to the traditional paraphrase evaluation for generation quality with reference-oriented BLEU [46] and perplexity (PPL) based on open-source GPT2-base model [47]; However, an ideal paraphrase must reconcile both generation quality and diversity. We adopt the distinct n-gram [48, div-4] for diversity, which measures the ratio of unique n-gram tokens to the total generations. Considering the stochastic nature of variational and diffusion generation, we have gathered the top-5 beam search samples for deterministic baselines (including fine-tuned LLMs), as well as 5 random samples for both the variational baselines and our proposed method with or without the controller. Quality metrics are averaged for fair comparison.
Note that reference-oriented metrics contradict the evaluation of generation diversity, where the paraphrase can sacrifice either quality for diversity via nonsense generation, or diversity via duplication. Therefore, we adopt the iBScore [49] as the primary metric, which measures the semantic via embedding similarity, namely the BERTScore [50], and punishes source duplication by source BLEU:
@@
where we calculate BERTScore via deBERTa-large [51].
Lastly, we additionally refer to the human evaluation (HE) in work [52] and sample 100 paraphrases from each test set for human evaluation ranging from 00 to $6$ by the following criterion:
  • 0: The paraphrase is nonsensical, failing to convey coherent meaning or with little expressive changes.
  • 1: The paraphrase is nonsensical, failing to convey coherent meaning or with unfluent expression.
  • 2: There is a certain amount of expression changes in the paraphrase, or significant changes that partially preserve the meaning of the source text, albeit with some unfluency or omissions.
  • 3: There is a certain amount of expression changes in the paraphrase with better fluency, albeit some inaccuracies.
  • 4: There are expressive changes in paraphrase, but it largely retains its semantics with minor issues such as slight grammar errors.
  • 5: There are expressive changes in paraphrase while retaining the semantics with fluency.
  • 6: There are creative or exceeding expressions in paraphrases while accurately conveying the full meaning of the source text without errors.
The scores support the half value in between. The sampled paraphrases are double-blind for random assignment, each with at least two scorers in case of bias. The above metrics are collected on generated samples and averaged for the final results.
HE
Transformer-base 3.9
BART-FT 4.465
T5-GPVAE 4.355
BART-CVAE 4.255
DiffuSeq 3.655
SeqDiffuSeq 3.32
Llama2-7b 4.63
Llama3.1-8b 4.915
LDP(ours) 4.585
w/ SG 4.75
Table 3: Results of human evaluation (HE). The top result is bolded, and the second best is underlined. LDP exceeds traditional and diffusion baselines. Compared to the SOTA LLMs, LDP achieves comparable results with the adaptor for semantics guidance (with SG).
@@ 4.2 Main Results
Table 1 presents the main results of our experiments, with human evaluation in Table 3. LDP achieves comparable or even superior performance compared to traditional baselines. Moreover, it outperforms other diffusion counterparts on QQP and Twitter test sets. LDP achieves the best performance amongst all baselines for the QQP test set.
Specifically, LDP improves overall results (iBScore), especially diversity (div-4) compared to traditional end-to-end paradigms such as fine-tuned BART, where we consider the diffusion module significantly improves the generation diversity. On the other hand, LDP also outperforms SOTA diffusion baselines such as SeqDiffuSeq in terms of BLEU and perplexity, where we consider the LDP to implement a better bridge between the diffusion process and the discrete texts than rounding. Additionally, unlike the intuition to introduce external features, the original case output improves greatly when we guide the generation semantics with a segment of source inputs as shown in Table 2. Though fine-tuned LLM outperforms others in human evaluations, it requires significantly larger resources and a pre-training corpus with significantly weaker reference-oriented BLEU scores. In contrast, LDP (with or without semantic guidance) achieves comparable results with substantially reduced overhead.
As shown in Table 4, we generate several paraphrases by different latent sampling for generations. Though DiffuSeq and BART-CVAE are SOTA generators for diversity, they still yield resemble paraphrases. LDP, on the other hand, yields better and more diverse paraphrases.
@@
Dinoiser[15] 26.07 -
Diff-Glat[56] 29.86 -
SeqDiffuSeq[41] 24.32 -
DiffuSeq[40] 24.13 83.65
@@
We additionally include more recent text diffusion generators for comparison for QQP tests. Table 5 shows that LDP achieves SOTA BLEU and BERTScore amongst the diffusion baselines.
Overall, LDP with semantic controller achieves the best iBScore amongst traditional and diffusion baselines, indicating a high-quality and diverse paraphrase generation.
@@ 5.1 Inference Efficiency
Lapse (s)
Transformer 235 (206.3$\times$)
BART-FT 238 (203.7$\times$)
T5-GPVAE 3909 (12.4$\times$)
BART-CVAE 252 (192.4$\times$)
DiffuSeq(DDIM 2000) 48480 (1$\times$)
DiffuSeq(DDIM 500) 11530 (4.2$\times$)
SeqDiffuSeq(DDIM 2000) 13851 (3.5$\times$)
LDP(ours) 290 (167.2$\times$)
Table 6: Inference overhead (seconds) with acceleration rate (in brackets) on the QQP validation set, where we set the DiffuSeq as the efficiency baseline.
By eliminating the rounding process in text diffusion models, LDP achieves improved efficiency. We compare several baseline efficiencies under their best-generation performance. As shown in Table 6, LDP is $167.2$ times faster than DiffuSeq and $50$ times faster than SeqDiffuSeq, where the diffusion overhead takes up only $18\%$ of the total timelapse. Consequently, our approach achieves a similar efficiency compared to that of the autoregressive baselines with only enc-dec overheads.
@@ 5.2 Domain Adaptation by Controller
adapted 36.46 14.65 (+4.49)
Table 7: Domain adaptation by controller. The controller retains the diffusion adaptation
The LDP with a controller retains its adapter capabilities, making it equally suitable for domain adaptation. We additionally adopt ChatGPT-augmented paraphrase dataset [57, ChatGPT-Aug] as the novel data domain, then sample $10$k sentences as the test set. ChatGPT-Aug dataset consists of paraphrases generated by ChatGPT from an ensembled dataset including QQP, SQuAD 2.0 [58] and CNN news [59].
We first train the LDP on QQP data for $200$k steps, then fine-tune it with a controller for ChatGPT-Aug data for $40$k steps. The controller fine-tuning follows the same routine where the ChatGPT-Aug data pairs are training sources and references, with additional keywords from its reference as controller inputs. Table 7 shows that adaptation by the controller is viable, where the fine-tuned controller substantially improves ChatGPT-Aug test BLEU, with only minor loss for the original test domain. Note that, the performance of the original data domain is retained by inference without the controller, which outstands from traditional data adaptation by fine-tuning.
@@ 5.3 Ablation Study for Samplers
The diffusion sampler plays an important role during inference, thus we conduct an ablation study on the diffusion sampler for comparison. Intuitively, the diffusion generation improves by more steps given the same sampler. Due to the truncation errors introduced by rounding, ODE-based samplers are not apt for text diffusion with rounding, such as DiffuSeq and SeqDiffuSeq. Thus, we additionally adopt DDIM sampler [12] with 500 and 200 sampling steps, which is also adopted for diffusion baselines. Table 8 shows that LDP still outperforms the DiffuSeq and SeqDiffuSeq given the same sampler setting. Notably, LDP still outperforms the baselines by fewer sampling steps, which were expected to be inferior.
@@ 5.4 Question Generation Ability
Considering the compatibility of LDP besides paraphrasing, we additionally validate question generation, which aims to generate the exact question by given issue descriptions. Question generation focuses on semantics upon relatively flexible utterance, thus we evaluate the BLEU and BERTScore on Quasar-T tests [60].
As shown in Table 9, LDP with enforced semantics also achieves a comparable BLEU score with SOTA, TESS [55], with much improved BERTScore.
@@ 6 Related Work
Paraphrase generation is commonly modeled by end-to-end sequence generation, such as fine-tuned Transformers [5, 31, 20]. However, such deterministic generation fails to ensure diversity, thus some researchers [61, 19, 62, 63] introduce external features to amend. Others turn to the variational generation for diversity [23]. To amend the inferior generation by latent representation, some researchers further introduce pre-trained text encoder and decoder for generation [9, 39].
On the other hand, the diffusion model has been a popular Markov variational generation in recent years. Existing text diffusions cater to the discrete nature of language for versatile generations. DiffuSeq [40] employs a single Transformer encoder and partial noising process to extend Diffusion-LM [13] for end-to-end text generations. They introduce discrete sampling for diffusion intervals called ‘rounding’, to ensure generation quality. However, rounding introduces truncation errors that hinder efficient skip-step diffusion sampler. BG-Diffuseq [54] aims to narrow the gap between training and sampling for text diffusion via incorporating distance penalty and adaptive decay sampling. Dinoiser [15], turns to mitigate truncation errors by manipulating the noise scale scheduled for text diffusion training and inference to improve its efficacy. On the other hand, SeqDiffuSeq [41] turns to a continuous diffusion modeling within an on-the-fly enc-dec framework. Similarly, DiffuSIA [53] introduces an enc-dec diffusion via spiral interaction. Diff-GLAT [56] incorporates residual glancing sampling, a text reconstruction via enc-dec, with its dropout rate as the noise scale. TESS [55] proposes a logit simplex space rather than embedding space for text diffusion.
@@ 7 Conclusion
We propose a novel paraphrase generation via the controllable latent diffusion model, LDP, which can further incorporate flexible semantic controls to guide the generation, e.g., only input segments rather than external features. Experiments show that LDP generates better paraphrases with superior efficiency than its diffusion counterparts. Further analysis shows that LDP largely retains the strengths of diffusion generation and is versatile enough for similar text generation methods, such as question generation and domain adaptation.
Overall, LDP strikes a better balance between generation quality and diversity than mainstream baselines.
{competinginterest}
@@
{acknowledgement}
This work is supported by the National Science Foundation of China (No. 62376116, 62176120), the Liaoning Provincial Research Foundation for Basic Research (No. 2022-KF-26-02), a research project of Nanjing University-China Mobile Joint Institute.